# ArMeme: Propagandistic Content in Arabic Memes

### v1

@@ Abstract
With the rise of digital communication memes have become a significant medium for cultural and political expression that is often used to mislead audience. Identification of such misleading and persuasive multimodal content become more important among various stakeholders, including social media platforms, policymakers, and the broader society as they often cause harm to the individuals, organizations and/or society. While there has been effort to develop AI based automatic system for resource rich languages (e.g., English), it is relatively little to none for medium to low resource languages. In this study, we focused on developing an Arabic memes dataset with manual annotations of propagandistic content. [^1] Propaganda is a form of communication designed to influence people’s opinions or actions toward a specific goal, employing well-defined rhetorical and psychological techniquesfor Propaganda Analysis (1938). We annotated $\sim 6K$ Arabic memes collected from various social media platforms, which is a first resource for Arabic multimodal research. We provide a comprehensive analysis aiming to develop computational tools for their detection. We will make them publicly available for the community.
@@
Firoj Alam1, Abul Hasnat2,3, Fatema Ahmed1, Md Arid Hasan4, Maram Hasanain1 1Qatar Computing Research Institute, HBKU, Qatar 2Blackbird.AI, USA, 3APAVI.AI, France 4University of New Brunswick, Fredericton, NB, Canada {fialam,mhasanain,fakter}@hbku.edu.qa, mhasnat@gmail.com, arid.hasan@unb.ca
@@ 1 Introduction
Social media platforms have enabled people to post and share content online. A significant portion of this content provides valuable resources for initiatives such as citizen journalism, raising public awareness, and supporting political campaigns. However, a considerable amount is posted and shared to mislead social media users and to achieve social, economic, or political agendas. In addition, the freedom to post and share content online has facilitated negative uses, leading to an increase in online hostility, as evidenced by the spread of disinformation, hate speech, propaganda, and cyberbullying Brooke (2019); Joksimovic et al. (2019); Schmidt and Wiegand (2017); Davidson et al. (2017); Da San Martino et al. (2019a); Van Hee et al. (2015). A lack of media literacy [^2] Media literacy encompasses the ability to access, analyze, evaluate, and create media in various forms. is also a major factor contributing to the spread of misleading information on social media Zannu et al. (2024). This can lead to the uncritical acceptance and sharing of false or misleading content, which can quickly disseminate through social networks. In their study, Zannu et al. (2024) highlight the crucial role of media literacy in mitigating the spread of fake news among users of platforms such as Instagram and Twitter.
Figure 1: Examples of Arabic memes representing different categories.
Online content typically consists of different modalities, including text, images, and videos. Disinformation, misinformation, propaganda, and other harmful content are shared across all these modalities. Recently, the use of Internet memes have become very popular on these platforms. A meme is defined as “a collection of digital items that share common characteristics in content, form, or stance, which are created through association and widely circulated, imitated, or transformed over the Internet by numerous users” Shifman (2013). Memes typically consist of one or more images accompanied by textual content Shifman (2013); Suryawanshi et al. (2020). While memes are primarily intended for humor, they can also convey persuasive narratives or content that may mislead audiences. To automatically identify such content, research efforts have focused on addressing offensive material Gandhi et al. (2020), identifying hate speech across different modalities Gomez et al. (2020); Wu and Bhandary (2020), and detecting propaganda techniques in memes Dimitrov et al. (2021a).
@@
Among the various types of misleading and harmful content, the spread of propagandistic content can significantly distort public perception and hinder informed decision-making. To address this challenge, research efforts have been specifically directed towards defining techniques and tackling the issue in different types of content, including news articles Da San Martino et al. (2019), tweets Alam et al. (2022b), memes Dimitrov et al. (2021a), and textual content in multiple languages Piskorski et al. (2023a). Most of these efforts have focused on English, with relatively little attention given to Arabic. Prior research on Arabic textual content includes studies presented at WANLP-2022 and ArabicNLP-2023 Alam et al. (2022b); Hasanain et al. (2023). However, for multimodal content, specifically memes, there are no available datasets or resources. To address this gap, we have collected and annotated a dataset consisting of approximately 6,000 memes, categorizing them into four categories (as shown in Figure 1) to identify propagandistic content. Below we briefly summarize the contribution of our work.
@@
  • Releasing the dataset to the community. [^3] Please contact the authors for early access to the dataset. The dataset and annotation guideline will be beneficial for research to develop automatic systems and enhance media literacy.
@@ 2 Related Work
The widespread use of social media has become one of the main ways of sharing information and is also responsible for creating and spreading misinformation and propaganda among users. Propagandistic techniques often utilize various types of content, such as fake news and doctored images, across multiple media platforms, frequently employing tools like bots. This information is distributed in diverse forms, including textual, visual, and multi-modal. To mitigate the impact of propaganda in online media, researchers have been developing resources and tools to identify and debunk such content.
@@ 2.1 Persuasion Techniques Detection
Early research on propaganda identification relies on the entire document to identify whether the content is propaganda, while recent studies focus on social media content Dimitrov et al. (2021b), news articles Da San Martino et al. (2019b), political speech Partington and Taylor (2017), arguments Habernal et al. (2017, 2018), and multimodal content Dimitrov et al. (2021a). Barrón-Cedeno et al. (2019) developed a binary classification (propaganda and non-propaganda) corpus to explore writing style and readability levels. An alternative approach followed by Habernal et al. (2017, 2018) to identify persuasion techniques within the texts constructing a corpus on arguments. Moreover, the study of Da San Martino et al. (2019b) developed a span-level propaganda detection corpus from news articles and annotated in eighteen propaganda techniques.
Piskorski et al. (2023b) developed a dataset from online news articles into twenty-two persuasion techniques containing nine languages to address the multilingual research gap. Following the previous work, Piskorski et al. (2023a) and SemEval-2024 task 4 focus on resource development to facilitate the detection of multilingual persuasion techniques. Focusing on multimodal persuasion techniques for memes, Dimitrov et al. (2021a) created a corpus containing 950 memes and investigated pretrained models for both unimodal and multimodal memes. The study of Chen et al. (2024) proposed a multimodal visual-textual object graph attention network to detect persuasion techniques from multimodal content using the dataset described in Piskorski et al. (2023b). In a recent shared task, Dimitrov et al. (2024) introduced a multilingual and multimodal propaganda detection task, which attracted many participants. The participants’ systems included various models based on transformers, CNNs, and LLMs.
@@ 2.2 Multimodal Content
The study of multimodal content has gained popularity among researchers for propaganda detection due to the effectiveness of multimodal content in spreading propaganda information and creating positive impacts among the targeted audience. Sharma et al. (2022) presented propaganda can be used to cause several types of harm including hate, violence, exploitation, etc. while spreading mis- and dis-information is also one of the main reasons Alam et al. (2022a). The study of Volkova et al. (2019) presented an in-depth analysis of multimodal content for predicting misleading information from news. Additionally, the deception and disinformation analysis on social media platforms using multimodal content in multilingual settings has been studied by Glenski et al. (2019). Moreover, hateful memes Kiela et al. (2020), propaganda in visual content Seo (2014), emotions and propaganda Abd Kadir et al. (2016) also studied by the researchers in the past few years.
Recent studies focusing on fine-tuning visual transformer models such as ViLBERT Lu et al. (2019), Multimodal Bitransformers Kiela et al. (2019), and VisualBERT Li et al. (2019). Cao et al. (2022) study focuses on multimodal hateful meme identification using prompting strategies by adopting Prakash et al. (2023). Hee et al. (2024) studied hate speech content moderation and discussed recent advancements leveraging large models.
Compared to previous studies, our work differs in that we provide the first resource for Arabic. Additionally, our annotation guidelines and data collection procedures for memes may be useful for other languages.
@@ 3 Dataset
Our data collection process involve several steps as highlighted in the Figure 2. We manually selected public groups from Facebook, Instagram, and Pinterest. In addition, we have also collected memes from Twitter using a set of keywords as listed in the Figure  3 (in Appendix). Our data curation consists of a series of steps as discussed below.
Manual selection of groups, links and keywords:
Focusing on the mentioned sources we have manually selected public groups, which contains post on public figures, celebrity, and mentions about politics.
@@
Given that Facebook, Instagram and Pinterest do provide API or do not allow automatic crawling images, therefore, we developed a semi-automatic approach to crawl images from these platforms. The steps include manually loading images and then crawl the images that are loaded on the browser. For the Twitter (X-platform), we used the keywords to crawl tweets, which consists of media/image.
@@
We used EasyOCR [^4] https://github.com/JaidedAI/EasyOCR to extract text from memes. Memes with no extracted OCR text were filtered out.
@@
For the manual annotation, we first prepared an annotation guideline to assist the annotators. The annotation tasks consist of two phases: (1) Phase 1 (meme categorization): labeling memes as (i) not-meme, (ii) other, (iii) not propaganda, or (iv) propaganda; (2) Phase 2 (text editing): editing the text to fix OCR errors. To facilitate the annotation tasks, we developed an annotation platform as presented in Appendix D. The details of the annotation guidelines are reported in Appendix C. Note that we developed the annotation guidelines in English, which were then translated into the Arabic language as detailed in Section (Appendix Section E). In phase 1 (meme categorization), each meme was annotated by three annotators and final label is decided based on majority agreement.
In Figure 1, we provide examples of memes representing different categories. Figure 1(a) depicts a couple in what appears to be a therapy session. The therapist asks, “Do you feel your wife is controlling you?” The wife responds, “No, I don’t feel so.” It is evident that the question was directed towards the husband, yet the wife answers instead of him. The irony lies in her controlling the conversation when her control is the subject of discussion. This meme attempts to humorously portray the stereotypical notion that wives are controlling in marriages. Figure 1(b) employs a play on words to create humor but does not contain any propagandistic techniques. Figure 1(c) features a meme that uses an image of a scene with dialogue and added text to create humor. However, it was categorized as “other” because the dialogues were in English, rather than “not propagandistic” or “propagandistic.” Figure 1(d) shows a picture of book covers, which might have been part of an advertisement.
@@
For the Phase 1 annotation, we computed annotation agreement using various evaluation measures, including Fleiss’ kappa, Krippendorff’s alpha, average observed agreement, and majority agreement. The resulting scores were 0.529, 0.528, 0.755, and 0.873, respectively. Based on the value of Krippendorff’s alpha, we can conclude that our annotation agreement score indicates moderate agreement. [^5] Note that Kappa values of 0.21–0.40, 0.41–0.60, 0.61–0.80, and 0.81–1.0 correspond to fair, moderate, substantial, and perfect agreement, respectively Landis and Koch (1977). In the final label selection, we excluded the $\sim$200 memes on which the annotators disagreed. In the second phase, we mainly edited text to fix the OCR errors, which has been done by a single annotator. To ensure the quality of the editing phase a random samples were checked by an expert annotator and periodically provided feedback. Note that the post-editing has been done for only propagandistic and not-propagandistic memes. It is to reduce the cost of the annotation and to further annotate them with span-level propaganda technique annotation.
Statistics:
Table 1 shows the number of memes for each category. For the rest of the experiments, the data was split into train, dev, and test as shown in the table. The dataset comprises a total of 5,725 annotated samples, with “Not propaganda” covers over half of the dataset ($\sim$66%), followed by “Propaganda.” The “Not-meme“ and “Other“ classes are significantly smaller in comparison. The distribution indicates a significant class imbalance, particularly between “Not propaganda” and the other classes, which could affect model training and performance.
In Table 2, we report the distribution of the dataset across different sources. The annotated number of memes reflects the memes we collected from various sources, as detailed in Table 5. We have the highest number of memes collected and annotated from Instagram. A very small number from Twitter is due to different image filtering steps. As shown in Table 2 the prevalence of propagandistic memes is relatively higher on Facebook than that of non-propagandistic memes.
@@ 4.2.2 Small Language Models (SLMs)
For the text-based unimodal model, we transformed text into $n$-gram (n=1) format using a tf-idf representation, considering the top 5,000 tokens, and trained an SVM model with a parameter value of $C=1$. Additionally, we fine-tuned several pre-trained transformer models (PLMs). These included the monolingual transformer model AraBERT (Antoun et al., 2020), Qarib Abdelali et al. (2021) and multilingual transformers such as multilingual BERT (mBERT) Devlin et al. (2019), and XLM-RoBERTa (XLM-r) (Conneau et al., 2019). We used the Transformer toolkit Wolf et al. (2019) for the experiment. Following the guidelines outlined in (Devlin et al., 2019), we fine-tuned each model using the default settings over three epochs. Due to instability, we performed ten reruns for each experiment using different random seeds, and we picked the model that performed best on the development set. We provided the details of the parameters settings in Appendix B.

### v2

@@ Abstract
[…]
utf8
@@
WARNING: This paper contains examples which may be disturbing to the reader
Firoj Alam1, Abul Hasnat2,3, Fatema Ahmad1, Md Arid Hasan4, Maram Hasanain1 1Qatar Computing Research Institute, HBKU, Qatar 2Blackbird.AI, USA, 3APAVI.AI, France 4University of New Brunswick, Fredericton, NB, Canada {fialam,fakter,mhasanain}@hbku.edu.qa, mhasnat@gmail.com, arid.hasan@unb.ca
[…]
utf8
ArMeme: Propagandistic Content in Arabic Memes
WARNING: This paper contains examples which may be disturbing to the reader
Firoj Alam1, Abul Hasnat2,3, Fatema Ahmad1, Md Arid Hasan4, Maram Hasanain1 1Qatar Computing Research Institute, HBKU, Qatar 2Blackbird.AI, USA, 3APAVI.AI, France 4University of New Brunswick, Fredericton, NB, Canada {fialam,fakter,mhasanain}@hbku.edu.qa, mhasnat@gmail.com, arid.hasan@unb.ca
@@ 1 Introduction
Social media platforms have enabled people to post and share content online. A significant portion of this content provides valuable resources for initiatives such as citizen journalism, raising public awareness, and supporting political campaigns. However, a considerable amount is posted and shared to mislead social media users and to achieve social, economic, or political agendas. In addition, the freedom to post and share content online has facilitated negative uses, leading to an increase in online hostility, as evidenced by the spread of disinformation, hate speech, propaganda, and cyberbullying Brooke (2019); Joksimovic et al. (2019); Schmidt and Wiegand (2017); Davidson et al. (2017); Da San Martino et al. (2019); Van Hee et al. (2015). A lack of media literacy [^2] Media literacy encompasses the ability to access, analyze, evaluate, and create media in various forms. is also a major factor contributing to the spread of misleading information on social media Zannu et al. (2024). This can lead to the uncritical acceptance and sharing of false or misleading content, which can quickly disseminate through social networks. In their study, Zannu et al. (2024) highlight the crucial role of media literacy in mitigating the spread of fake news among users of platforms such as Instagram and Twitter.
Figure 1: Examples of images representing different categories.
@@
Online content typically consists of different modalities, including text, images, and videos. Disinformation, misinformation, propaganda, and other harmful content are shared across all these modalities. Recently, the use of Internet memes have become very popular on these platforms. A meme is defined as “a collection of digital items that share common characteristics in content, form, or stance, which are created through association and widely circulated, imitated, or transformed over the Internet by numerous users” Shifman (2013). Memes typically consist of one or more images accompanied by textual content Shifman (2013); Suryawanshi et al. (2020). While memes are primarily intended for humor, they can also convey persuasive narratives or content that may mislead audiences. To automatically identify such content, research efforts have focused on addressing offensive material Gandhi et al. (2020), identifying hate speech across different modalities Gomez et al. (2020); Wu and Bhandary (2020), and detecting propaganda techniques in memes Dimitrov et al. (2021a). Among the various types of misleading and harmful content, the spread of propagandistic content can significantly distort public perception and hinder informed decision-making. To address this challenge, research efforts have been specifically directed towards defining techniques and tackling the issue in different types of content, including news articles Da San Martino et al. (2019), tweets Alam et al. (2022b), memes Dimitrov et al. (2021a), and textual content in multiple languages Piskorski et al. (2023a). Most of these efforts have focused on English, with relatively little attention given to Arabic. Prior research on Arabic textual content includes studies presented at WANLP-2022 and ArabicNLP-2023 Alam et al. (2022b); Hasanain et al. (2023). However, for multimodal content, specifically memes, there are no available datasets or resources. To address this gap, we have collected and annotated a dataset consisting of approximately 6,000 memes, categorizing them into four categories (as shown in Figure 1) to identify propagandistic content. Below we briefly summarize the contribution of our work.
@@
  • Releasing the dataset to the community. [^3] Dataset is released under CC-BY-NC-SA through https://huggingface.co/datasets/QCRI/ArMeme. The dataset and annotation guideline will be beneficial for research to develop automatic systems and enhance media literacy.
@@ 2 Related Work
Social media has become one of the main ways of sharing information. Its widespread use and reach is also responsible for creating and spreading misinformation and propaganda among users. Propagandistic techniques can be found in various types of content, such as fake news and doctored images, across multiple media platforms, frequently employing tools like bots. Furthermore, such information is distributed in diverse forms, including textual, visual, and multi-modal. To mitigate the impact of propaganda in online media, researchers have been developing resources and tools to identify and debunk such content.
@@ 2.1 Persuasion Techniques Detection
Early research on propaganda identification relies on the entire document to identify whether the content is propaganda, while recent studies focus on social media content Dimitrov et al. (2021b), news articles Da San Martino et al. (2019), political speech Partington and Taylor (2017), arguments Habernal et al. (2017, 2018), and multimodal content Dimitrov et al. (2021a). Barrón-Cedeno et al. (2019) developed a binary classification (propaganda and non-propaganda) corpus to explore writing style and readability levels. An alternative approach followed by Habernal et al. (2017, 2018) to identify persuasion techniques within texts constructing a corpus on arguments. Moreover, Da San Martino et al. (2019) developed a span-level propaganda detection corpus from news articles and annotated in eighteen propaganda techniques. Piskorski et al. (2023b) developed a dataset from online news articles into twenty-two persuasion techniques containing nine languages to address the multilingual research gap. Following the previous work, Piskorski et al. (2023a) and SemEval-2024 task 4 focus on resource development to facilitate the detection of multilingual persuasion techniques. Focusing on multimodal persuasion techniques for memes, Dimitrov et al. (2021a) created a corpus containing 950 memes and investigated pretrained models for both unimodal and multimodal memes. The study of Chen et al. (2024) proposed a multimodal visual-textual object graph attention network to detect persuasion techniques from multimodal content using the dataset described in Piskorski et al. (2023b). In a recent shared task, Dimitrov et al. (2024) introduced a multilingual and multimodal propaganda detection task, which attracted many participants. The participants’ systems included various models based on transformers, CNNs, and LLMs.
@@ 2.2 Multimodal Content
The study of multimodal content has gained popularity among researchers for propaganda detection due to its effectiveness of in spreading propagandastic information and creating impact among the targeted audience. Sharma et al. (2022) presented that propaganda can be used to cause several types of harm including spreading hate, violence, exploitation, etc., while spreading mis- and dis-information is also one of the main reasons Alam et al. (2022a). The study of Volkova et al. (2019) presented an in-depth analysis of multimodal content for predicting misleading information from news. Additionally, the deception and disinformation analysis on social media platforms using multimodal content in multilingual settings has been studied by Glenski et al. (2019). Moreover, hateful memes Kiela et al. (2020), propaganda in visual content Seo (2014), emotions and propaganda Abd Kadir et al. (2016) also studied by the researchers in the past few years. Recent studies focusing on fine-tuning visual transformer models such as ViLBERT Lu et al. (2019), Multimodal Bitransformers Kiela et al. (2019), and VisualBERT Li et al. (2019). Cao et al. (2022) study focuses on multimodal hateful meme identification using prompting strategies by adopting Prakash et al. (2023). Hee et al. (2024) studied hate speech content moderation and discussed recent advancements leveraging large models. Compared to previous studies, our work differs in that we provide the first resource for Arabic. Additionally, our annotation guidelines and data collection procedures for memes may be useful for other languages.
@@ 3 Dataset
3.1 Data Collection
Our data collection process involves several steps as highlighted in the Figure 2. We manually selected public groups and contents from Facebook, Instagram, and Pinterest. In addition, we have also collected memes from Twitter using a set of keywords as listed in the Figure  3. Our data curation consists of a series of steps as discussed below.
Figure 3: Keywords used to collect tweets.
Manual selection of groups, links, and keywords:
Focusing on the mentioned sources, we have manually selected public groups, which contains post on public figures, celebrity, and discussions about politics. In Table 1, we provide the sources of the dataset, number of groups, and number of images we have collected.
Source # of Group # of Images
Facebook 19 5,453
Instagram 22 107,307
Pinterest - 11,369
Twitter - 5,369
Total 129,498
Table 1: Statistics of the initial data collection.
@@ 3.1 Data Collection
Given that Facebook, Instagram, and Pinterest do not provide API or do not allow automatic crawling images, therefore, we developed a semi-automatic approach to crawl images from these platforms. The steps include manually loading images and then crawl the images that are loaded on the browser. For the Twitter (X-platform), we used the keywords to crawl tweets, which consists of media/image.
3.2 Filtering
@@
We used EasyOCR [^4] https://github.com/JaidedAI/EasyOCR to extract text from memes. Memes with no extracted or detected text were filtered out. The reasons for choosing EasyOCR are: (a) it is a ready-to-use OCR with 80+ supported languages and all popular writing scripts, and (b) it includes the implementation of the state-of-the-art, highly efficient real-time scene text detection module Liao et al. (2022), called DBnet, which uses differentiable binarization and adaptive scale fusion.
@@ 3.3 Annotation
For the manual annotation, we first prepared an annotation guideline to assist the annotators. To facilitate the annotation tasks, we developed an annotation platform as presented in Appendix C. The details of the annotation guidelines are reported in Appendix B. Note that we developed the annotation guidelines in English, (see Section B), which were then translated into the Arabic language. Translating the guideline in native language was indeed important and also inspired by prior work Alam et al. (2021); Hasanain et al. (2024a). The idea is not only to make the annotation task more convenient but also capture different linguistic aspects. The guidelines included several examples of memes. It was reviewed by several NLP experts who are also native Arabic speakers. In Figure 1, we provide examples of images and memes representing different categories. Figure 1(a) depicts a couple in what appears to be a couples therapy session. The therapist asks the husband, “Do you feel your wife is controlling you?” The wife responds, “No, I don’t feel so.” It is evident that the question was directed towards the husband, yet the wife answers instead of him. The irony lies in her controlling the conversation when her control is the subject of discussion. This meme attempts to humorously portray the stereotypical notion that wives are controlling in marriage. Figure 1(b) employs a play on words to create humor. The Arabic word that means "different" is similar to the Arabic word "retarded", except in the position of two letters. However, this meme does not contain any propagandistic techniques. Figure 1(c) features a meme that uses an image of a scene from TV with dialogues and added text to create humor. However, it was categorized as “other” because the dialogues were in English, rather than “not propagandistic” or “propagandistic”. Figure 1(d) shows a picture of book covers, which might have been part of an advertisement. This image was labelled as "not meme". The annotation tasks consist of two phases:
  • Phase 1 (image categorization): labeling images shown on the platform as (i) not-meme, (ii) other, (iii) not propaganda, or (iv) propaganda. Each image was annotated by three annotators and the final label is decided based on majority agreement.
  • Phase 2 (text editing): editing the text to fix OCR errors. This step was performed only for images labelled as meme, and propagandistic or not propagandistic. Further details is mentioned in B.2.
Annotation Team: The team in phase 1 consisted of three members, and in phase 2, consisted of one member. All annotators are native Arabic speakers holding at least a bachelor’s degree. Our in-house expert annotator provided them with several iterations of training, supervised and monitored their work, and handled quality control throughout the entire annotation process. This quality assurance included periodic checks of random annotation samples and providing feedback. Since the institute requires the signing of a Non-Disclosure Agreement (NDA), each annotator signed an NDA after being made aware of the institute’s terms and conditions. They were compensated at the same rate as charged by external companies. Annotation platform: We utilized our in-house annotation platform for the annotation task. Separate annotation interfaces were designed for each phase.
@@
For the Phase 1 annotation, we computed annotation agreement using various evaluation measures, including Fleiss’ kappa, Krippendorff’s alpha, average observed agreement, and majority agreement. The resulting scores were 0.529, 0.528, 0.755, and 0.873, respectively. Based on the value of Krippendorff’s alpha, we can conclude that our annotation agreement score indicates moderate agreement. [^5] Note that Kappa values of 0.21–0.40, 0.41–0.60, 0.61–0.80, and 0.81–1.0 correspond to fair, moderate, substantial, and perfect agreement, respectively Landis and Koch (1977). In the final label selection, we excluded the $\sim$200 memes on which the annotators disagreed. In the second phase, we mainly edited text to fix the OCR errors, which has been done by a single annotator. To ensure the quality of the editing phase, random samples were checked by an expert annotator and periodically provided feedback. Note that the post-editing has been done for only propagandistic and non-propagandistic memes. It is to reduce the cost of the annotation, and to further annotate them with span-level propaganda techniques.
3.4 Statistics
Table 2 shows the number of memes for each category. For the rest of the experiments, the data was split into train, dev, and test as shown in the table. The dataset comprises a total of 5,725 annotated samples, with “Not propaganda” covers over half of the dataset ($\sim$66%), followed by “Propaganda.” The “Not-meme“ and “Other“ classes are significantly smaller in comparison. The distribution indicates a significant class imbalance, particularly between “Not propaganda” and the other classes, which could affect model training and performance. In Table 3, we report the distribution of the dataset across different sources. The annotated number of memes reflects the memes we collected from various sources, as detailed in Table 1. We have the highest number of memes collected and annotated from Instagram. A very small number from Twitter is due to different image filtering steps. As shown in Table 3 the prevalence of propagandistic memes is relatively higher on Facebook than that of non-propagandistic memes.
@@ 4.2.2 Small Language Models (SLMs)
For the text-based unimodal model, we transformed text into $n$-gram (n=1) format using a tf-idf representation, considering the top 5,000 tokens, and trained an SVM model with a parameter value of $C=1$. Additionally, we fine-tuned several pre-trained transformer models (PLMs). These included the monolingual transformer model AraBERT (Antoun et al., 2020), Qarib Abdelali et al. (2021) and multilingual transformers such as multilingual BERT (mBERT) Devlin et al. (2019), and XLM-RoBERTa (XLM-r) (Conneau et al., 2019). We used the Transformer toolkit Wolf et al. (2019) for the experiment. Following the guidelines outlined in (Devlin et al., 2019), we fine-tuned each model using the default settings over three epochs. Due to instability, we performed ten reruns for each experiment using different random seeds, and we picked the model that performed best on the development set. We provided the details of the parameters settings in Appendix A.